Dot Product
Suppose that u and v are vectors in R to the seventh power both of length 2 the square root of 2 and that the length of u − v is also2 the square root of 2 . Then parallel to u plus v parallel-to = ____________ and the angle between u and v is _____________.
1
Expert's answer
2020-11-11T12:21:31-0500

|u-v|2 = |u|2 + |v|2 - 2*|u|*|v|*cos"\theta"

(2"\sqrt{2}")2 = (2"\sqrt{2}")2 + (2"\sqrt{2}")2 - 2*(2"\sqrt{2}")*(2"\sqrt{2}")*cos"\theta"

8 = 8 + 8 - 16*cos"\theta"

cos"\theta"=0.5

"\theta" = "\pi"/3

Suppose that u and v are vectors in R to the seventh power both of length 2 the square root of 2 and that the length of u − v is also2 the square root of 2 . Then parallel to u plus v parallel-to = k*(u+v) (where k=const) and the angle between u and v is "\pi"/3 radian.